OpenStudy (davidfedrick):

What is the common ratio of the sequence? -2, 6, -18, 54,...

OpenStudy (tkhunny):

Have you considered calculating the successive ratios? 6 / (-2) = ?? -18 / 6 = ?? 54 / (-18) = ?? etc...

OpenStudy (davidfedrick):

nah not yet.

OpenStudy (amorfide):

if we are talking about sequences and ratio, you have a geometric sequence the standard form of a geometric sequence is a, ar, ar^2, ar^3, ar^4.... ar^(n-1) so to get the common ratio, you must do an+1/an so let n=1 an=a1=a an+1=a2=ar ar/a=r let n=2 an=a2=ar an+1=a3=ar^2 ar^2/ar=r so in your sequence n=1, an=a1=-2 an+1=a2=6 6/-2=-3 n=2, an=a2=6 an+1=a3=-18 -18/6=-3 do this for all terms of your sequence to make sure they are the same r, ratio, throughout, hence this is your common ratio